The sum of three numbers is 88. The first number decreased by 5, the second number increased by 5, and the third number multiplied by 5 are all equal. Find the positive difference between the smallest and the largest of the three numbers.
a) 8 b) 35 c) 37 d) 45 e) none of these

Answer by richwmiller(17219)

a+b+c=88
a-5=b+5
a=b+10
b+5=c*5
c=(b+5)/5
a+b+c=88
b+10+b+((b+5)/5)=88
combine
2b+10+(b+5)/5=88
2b+(b+5)/5=78
multiply by 5
10b+b+5=78*5
11b=77*5
divide by 11
b=7*5
b=35
a=b+10
c=(b+5)/5
a=45
c=(35+5)/5
c=40/5=8
smallest =8
largest =45
45-8=37
c)37